### v2

# PINNfluence: Influence Functions for Physics-Informed Neural Networks

Abstract
Recently, physics-informed neural networks (PINNs) have emerged as a flexible and promising application of deep learning to partial differential equations in the physical sciences. While offering strong performance and competitive inference speeds on forward and inverse problems, their black-box nature limits interpretability, particularly regarding alignment with expected physical behavior. In the present work, we explore the application of influence functions (IFs) to validate and debug PINNs post-hoc. Specifically, we apply variations of IF-based indicators to gauge the influence of different types of collocation points on the prediction of PINNs applied to a 2D Navier-Stokes fluid flow problem. Our results demonstrate how IFs can be adapted to PINNs to reveal the potential for further studies. The code is publicly available at https://github.com/aleks-krasowski/PINNfluence.
1Fraunhofer HHI, 10587 Berlin, Germany
2Technische Universität Berlin, 10587 Berlin, Germany
3BIFOLD – Berlin Institute for the Foundations of Learning and Data, 10587 Berlin, Germany
[…]
$\ast$
equal contribution.
1 Introduction
Time and time again, deep learning approaches have proven themselves to be exceptionally capable at solving problems that were considered complicated and time-consuming, yielding fast inference times bundled with strong performance. Physics-informed neural networks (PINNs) [28] represent another recent iteration of that trend applied to the realm of partial differential equations (PDEs). By incorporating prior physical knowledge in the form of differential equations [7; 13; 14; 28], they admit a wide range of possible applications such as fluid mechanics, electromagnetics, disease modelling and optics, inter alia [2; 3; 5; 23]. Despite the power of PINNs, there remains the challenge of understanding and improving model behavior when things go wrong [19; 31; 32]. PINNs minimize complex composite losses, which complicates the tracing of poor performance to specific training points or conditions; this is the main motivation for this work.
Influence functions [16] (IFs) systematically assess the contribution of individual training points to a model’s behavior. Recent studies have shown that incorporating physical concepts like temporal causality improves PINN performance [10; 34], highlighting the role of underlying physical principles. By identifying key points, such as those near boundaries or critical regions, we can assess whether the PINN’s learned parameters align with the physics of the problem. For many PDEs, practitioners possess an intuitive understanding of the underlying physics. For example, in magnetostatics, the magnetic field is fundamentally generated by source terms. A well-trained PINN could be expected to reflect this relation. We aim to show how domain experts can use this approach to validate the model by identifying key training points, boundary conditions or physical principles that have the greatest impact on the model’s behavior. This also allows them to estimate the effect of inductive biases and observed data on the model’s predictions. Notably, PINNs also provide a powerful application domain for IFs and other data attribution methods, as models and datasets are usually much smaller than those used in computer vision or natural language processing. This makes the calculation of IFs computationally feasible and the effect of individual data points on the fit of the model is expected to be greater.
Our Contributions
  • We adapt influence functions to PINNs as a tool for improving their interpretability.
  • We derive heuristically motivated indicators based on influence functions to showcase how they can be leveraged by domain-knowledgeable users to test and validate PINNs.
  • We empirically demonstrate their usefulness in disambiguating between three different versions of PINNs applied to a Navier-Stokes problem.
2 Theoretical Background
Physics-Informed Neural Networks
[28] constitute a machine-learning-based approach to solving PDEs, which are ubiquitous in physical sciences. Let $\Omega\subseteq\mathbb{R}^{n}$ be an open domain and consider the initial boundary value problem (IBVP)
$\mathcal{N}[u](\boldsymbol{x})=0,\quad\boldsymbol{x}\in\Omega$ (1)
$\mathcal{B}[u](\boldsymbol{x})=0,\quad\boldsymbol{x}\in\partial\Omega\quad\text{.}$ (2)
Here, $\mathcal{N}$ and $\mathcal{B}$ are differential operators acting on the solution of the PDE $u:\Omega\to\mathbb{R}^{d}$, where $\boldsymbol{x}$ typically represents the spatial as well as possible temporal coordinates. Note that Eq. 2 is formulated to capture i.a. Dirichlet as well as von-Neumann conditions.
The goal of PINNs is to approximate the solution of a given differential equation using a neural network $\phi(\boldsymbol{x};\theta)$, which depends on parameters $\theta\in\Theta$ and an input $\boldsymbol{x}$. These parameters are optimized by training $\phi$ using a composite loss function, consisting of the PDE residual and boundary conditions $\mathcal{L}=\mathcal{L}_{\text{pde}}+\mathcal{L}_{\text{bc}}$, with $\mathcal{L}_{\text{pde}}=\frac{1}{N_{\text{pde}}}\sum_{i=1}^{N_{\text{pde}}}\left|\mathcal{N}[\phi(\boldsymbol{x}_{i};\theta)]\right|^{2}$, $\mathcal{L}_{\text{bc}}=\frac{1}{N_{\text{bc}}}\sum_{i=1}^{N_{\text{bc}}}\left|\mathcal{B}[\phi(\boldsymbol{x}_{i};\theta)]\right|^{2}$, and where $N_{\text{pde}}$ as well as $N_{\text{bc}}$ are the numbers of randomly sampled collocation from $\Omega$ and $\partial\Omega$, respectively. Note that one can also consider a data-driven regression loss, which we drop for the sake of simplicity.
Influence Functions
The question being addressed by influence functions [6; 11; 16] is how would the model’s behavior change if certain training points were not present in the training set? By linearly approximating the effect of leave-one-out retraining, IFs provide a possibility of studying a model through the lens of its training data. In order to adapt the influence functions to PINNs we will slightly generalize the approach of [16; 17]. We provide a full derivation and proof, including details on all necessary assumptions, in Section A.1.
Suppose we have trained a model by minimizing a loss $\frac{1}{N}\Sigma_{i}L(\boldsymbol{x}_{i};\theta)$ on a training data set $\mathcal{X}=\{\boldsymbol{x}_{1},\ldots,\boldsymbol{x}_{N}\}$ and derived an optimal parameter $\hat{\theta}\in\Theta$. Furthermore, consider a function $f:A\times\Theta\to\mathbb{R}$, which depends on the parameter $\theta\in\Theta$ and some other parameter $\boldsymbol{z}\in A$ (such as in [16], the loss $L$ for a single test point $\boldsymbol{z}$). Before discussing the case of removing and adding multiple training points, we begin with the case of adding a single training point to $\mathcal{X}$. Thus, we aim to approximate $f(\boldsymbol{z};\hat{\theta})-f(\boldsymbol{z};\hat{\theta}^{+})$ if $\hat{\theta}^{+}$ were optimized over an amended training data set that includes an additional training point $\boldsymbol{x}^{+}$. This can be approximated by
$\operatorname{Inf}_{f(\boldsymbol{z};\hat{\theta})}(\boldsymbol{x}^{+})\coloneq\nabla_{\theta}f(\boldsymbol{z};\hat{\theta})^{\top}\cdot\mathcal{H}_{\hat{\theta}}^{-1}\cdot\nabla_{\theta}L(\boldsymbol{x}^{+};\hat{\theta})\quad\text{,}$ (3)
where $\mathcal{H}_{\hat{\theta}}\coloneq\nabla^{2}_{\theta}\frac{1}{N}\sum_{i=1}^{N}L(\boldsymbol{x}_{i};\hat{\theta})$ denotes the Hessian w.r.t. $\theta$. This approach can be extended to the addition and removal of multiple training data points: Let $\mathcal{X}^{+}=\{\boldsymbol{x}_{1}^{+},\ldots,\boldsymbol{x}_{m^{+}}^{+}\}$ be the set of training points we want to add and $\mathcal{X}^{-}=\{\boldsymbol{x}_{1}^{-},\ldots,\boldsymbol{x}_{m^{-}}^{-}\}\subseteq\mathcal{X}$ the set we want to remove from our training set. The effect of retraining the model with added and removed training points can then be approximated by ${\operatorname{Inf}_{f(\boldsymbol{z};\hat{\theta})}(\mathcal{X}^{+},\mathcal{X}^{-})\coloneqq\sum_{i=1}^{m^{+}}\operatorname{Inf}_{f(\boldsymbol{z};\hat{\theta})}(\boldsymbol{x}_{i}^{+})-\sum_{i=1}^{m^{-}}\operatorname{Inf}_{f(\boldsymbol{z};\hat{\theta})}(\boldsymbol{x}_{i}^{-})}$. Hence, we can study the influence of single sampling points as well as regions on loss functions, any type of predictions or derived quantities. By studying the influence resulting therefrom, one can compare the model’s behavior with physical intuition. Using this method, we can also estimate the effect of adding and removing entire loss terms for composite losses, as well as several other types of manipulations (see Eqs. 15, 18, 21 and 23 in the appendix).
3 Experiments
Navier-Stokes Problem
The Navier-Stokes equations describe the motion of viscous Newtonian fluids and represent the cornerstone of fluid dynamics. Consequently, they are widely used in various applications. In the present work, we focus on a laminar, incompressible and time-independent fluid flow around a cylinder in a rectangular cavity with open lateral boundaries in 2D [30], see Fig. 1. The stationary equations of motions as well as the continuity equation are given as follows:
$(\boldsymbol{u}\cdot\nabla)\boldsymbol{u}+\frac{1}{\rho}\nabla p-\nu\nabla^{2}\boldsymbol{u}=0\ ,\qquad\operatorname{div}\boldsymbol{u}=0\ \text{,}$ (4)
where $\boldsymbol{u}(x,y)\in\mathbb{R}^{2}$ denotes the velocity field, $p(x,y)\in\mathbb{R}$ represents the pressure, $\rho$ is the density, and $\nu$ is the kinematic viscosity. The domain is given by $\Omega=(x_{\min},x_{\max})\times(y_{\min},y_{\max})\setminus C_{r}$, where $C_{r}$ denotes the cylinder with radius $r$. The boundary conditions at the bottom, top, and cylinder are given by $\boldsymbol{u}=0$. The inflow condition on the left has a parabolic profile.
Figure 1: Trained model ($\phi_{\text{good}}$) with predictions: $u_{1}$, $u_{2}$, $p$ (top to bottom).
For the rightward outflow condition, we use a directional "do nothing" condition [24]: $\nu\nabla u_{1}(x_{\text{max}},y)-(\frac{p}{\rho},0)^{\top}=0$. The used parameters of the geometry and PDE are given in A.2. The task of the PINN is then to predict $\boldsymbol{u}(x,y)$ and $p(x,y)$.
Experimental Setup
We train fully-connected PINNs consisting of $4$ layers with $64$ nodes each with GELU activation functions and $3$ outputs. These correspond to the solution of Eq. 4 that we want to predict: $u_{1}$, $u_{2}$ and $p$. For the training data set, we randomly draw $N_{\text{pde}}$ points inside the domain $\Omega$ and $N_{\text{bc}}$ from the boundary using Hammersley sampling. In essence, these are pairs of $(x,y)$ coordinates as we consider a 2D PDE. We train three models, (1) good model $\phi_{\text{good}}$ with the correct PDE and $N_{\text{pde}}=7500$ and $N_{\text{bc}}=2500$ that solves the task well compared to a reference solution [33]; (2) a broken model $\phi_{\text{broken}}$ with the same number of training points but with the $(u_{1}\cdot\nabla)u_{1}$ part of Eq. 4 missing, thus trained on wrong physics; (3) and a bad model $\phi_{\text{bad}}$ with the correct PDE but with $N_{\text{pde}}=1500$ and $N_{\text{bc}}=500$, such that it consequently performs badly. Each model is trained in two phases. First, Adam [15] is used as an optimizer for a total of $100$k steps, followed by further refining the model parameter’s using L-BFGS [21] for up to $25$k steps. We slightly adapted the DeepXDE library [22] to allow for more fine-grained gradient manipulation while using PyTorch [26] as our backend. For the test set, we draw a total of $36934$ points from $\Omega$ and $1288$ points from $\partial\Omega$ [33]. The influences are calculated using the NaiveInfluenceFunction from captum [18].
Empirical Indicators
Figure 2: Influence heatmap on the prediction of $u_{1}$ of a single training point ($\times$) close to the cylinder (bottom).
In this study, we aim to answer the following question using influence functions: Can we test whether the PINN has learned certain aspects of the underlying physical principles of the Navier-Stokes problem? Since predictions alone do not reveal the influence of training data or boundary conditions, we devise a set of heuristic indicators.
We propose two IF-based indicators that are designed to capture aspects of the physical processes inherent to the PDE (4) and to serve as exemplary metrics showcasing how expert users can design tests to validate PINNs against their prior knowledge of the problem. As a first indication metric, we concentrate on the flow direction of the Navier-Stokes problem: the inflow is given at the left boundary of the domain and the outflow lies on the opposite side. Figure 2 shows the influence pattern of a training point close to the cylinder.
We thus test a model on whether its influence patterns reflect this behavior. To define the directional indicator ($\operatorname{DI}$), which captures the relevance share associated with the collocation point $\boldsymbol{x}_{\text{test}}$ on the test points that follow in the direction of the flow (i.e. having larger $x$-component), we write:
$\operatorname{DI}(\boldsymbol{x}_{\text{test}}):=\frac{\sum_{\boldsymbol{x}_{\text{train}}:x_{\text{train}}>x_{\text{test}}}\left|\operatorname{Inf}_{f(\boldsymbol{x}_{\text{test}};\hat{\theta})}(\boldsymbol{x}_{\text{train}})\right|}{\sum_{\boldsymbol{x}_{\text{train}}}\left|\operatorname{Inf}_{f(\boldsymbol{x}_{\text{test}};\hat{\theta})}(\boldsymbol{x}_{\text{train}})\right|}\in[0,1]$ (5)
A larger value of $\operatorname{DI}$ implies that a given test point mainly influences the prediction further down in the direction of the flow rather than before and implies that the model has learned this underlying physical phenomenon of directed flows.
As a second indicator, we are interested in the fraction of influence that is associated with a given object in the input domain. In the present Navier-Stokes problem, the cylinder plays a central role in the flow field: it obstructs the flow.
(a) $\phi_{\text{good}}$
(b) $\phi_{\text{broken}}$
(c) $\phi_{\text{bad}}$
Figure 3: Average log values of $|\operatorname{Inf}_{\sum_{i}|\!|u(\boldsymbol{x}_{i};\theta)|\!|}(\boldsymbol{x})|$ over the domain $\overline{\Omega}$. The area of $C_{1.5r}$ is outlined in black.
Thus, the overall velocity and pressure fields should be heavily influenced by its position. To this end, we devise a region identifier, which is designed to capture the influence associated with a given object or region in the domain (in our case around the cylinder). For the region indicator ($\operatorname{RI}$), which sums up the relevance of removing a region $\Xi\subseteq\overline{\Omega}$, we write:
$\operatorname{RI}(\Xi):=\frac{\sum_{\boldsymbol{x}_{\text{train}}\in\Xi}\left|\operatorname{Inf}_{f(\boldsymbol{x}_{\text{test}};\hat{\theta})}(\boldsymbol{x}_{\text{train}})\right|}{\sum_{\boldsymbol{x}_{\text{train}}}\left|\operatorname{Inf}_{f(\boldsymbol{x}_{\text{test}};\hat{\theta})}(\boldsymbol{x}_{\text{train}})\right|}\in[0,1]$
A higher value of $\operatorname{RI}$ means that the region in question, such as the cylinder in our case, has an increased influence on the prediction for all other points. This is to be expected for a model that aims to accurately reflect the physics. In this investigation, we choose $\Xi=C_{1.5r}$, meaning that we remove the area corresponding to a circle with 1.5 times the radius of the cylinder, see Fig. 3.
Results
Table 1: Mean values for the Direction $\overline{\operatorname{DI}}$ and Region $\overline{\operatorname{RI}}$ Indicators, aggregated over all $\boldsymbol{x_{\text{test}}}$. A higher score indicates a closer alignment with the indicator’s assumptions.
$\overline{\operatorname{DI}}$ $\overline{\operatorname{RI}}(C_{1.5r})$
$f(x;\theta)$ $\phi_{\text{good}}$ $\phi_{\text{bad}}$ $\phi_{\text{broken}}$ $\phi_{\text{good}}$ $\phi_{\text{bad}}$ $\phi_{\text{broken}}$
$u_{1}$ 0.8 0.8 0.71 0.23 0.18 0.13
$u_{2}$ 0.82 0.77 0.69 0.26 0.17 0.14
$p$ 0.83 0.87 0.79 0.28 0.18 0.22
$|\!|\boldsymbol{u}|\!|$ 0.8 0.79 0.71 0.23 0.18 0.13
$\mathcal{L}_{\text{pde}}$ 0.79 0.76 0.73 0.21 0.14 0.16
$\mathcal{L}_{\text{bc}}$ 0.75 0.71 0.65 0.25 0.16 0.15
$\mathcal{L}$ 0.79 0.76 0.73 0.21 0.14 0.16
We evaluate the three different iterations of PINNs on the Navier-Stokes task. Figures 3(a) to 3(c) show the influence of individual training points on the whole test set, as heatmaps. For $\phi_{\text{good}}$, the most influential points are distributed close to the inflow, on the outflow and especially around the cylinder. This is expected, as these areas determine the fluid flow. Furthermore, we observe that for $\phi_{\text{bad}}$, the upper left boundary is overly influential. In contrast, $\phi_{\text{broken}}$’s influences are again distributed mostly around the cylinder. In Table 1, the aggregated mean values of the indicators are presented. Both $\phi_{\text{good}}$ and $\phi_{\text{bad}}$ show stronger $\operatorname{DI}$ performance. This suggests that both models have successfully learned to capture the dependence of the fluid flow on preceding points. When it comes to the $\operatorname{RI}$ indicator, this trend is even clearer. Here, the good model achieves superior performance, which suggests that the cylinder here has the biggest absolute influence, reflecting its importance for the fluid flow. It is important to note that these indicators are tools for investigating model behavior. Performing well on an indicator does not imply that the model is entirely correct, but rather that it does not fail in the specific aspect the indicator evaluates. Expert knowledge is essential for designing meaningful indicators and further investigation is needed to validate their robustness. Additionally, if the indicator’s assumptions are flawed, the results may be misleading and fail to provide sensible insights.
4 Conclusion
We showcased how influence functions can help to evaluate PINNs. In particular, we formulated heuristically motivated indicators designed to test certain physical concepts. By applying this method to a Navier-Stokes problem, we observed that the best-performing model consistently aligned with these indicators, confirming their utility in assessing model performance. Still, further exploration of IF-based markers is indicated to rigorously validate the broad applicability of the presented approach. Here, applying novel evaluation frameworks such as [1] could be helpful in future examinations. Using influence functions to improve PINN training through methods like resampling training points [20], enforcing temporal causality [34], or dynamically reweighing loss terms [4] – combined with IF-based training time techniques like TracIN [27] – presents a promising direction for future research. Additionally, further studies on the influence on parts of the loss as well as on different physical quantities, e.g. the vorticity, as well as applying the framework to different PDEs such as the heat equation, Maxwell equations and others appear to be logical next steps.
Acknowledgments
The authors want to thank Galip Ümit Yolcu for fruitful discussions. This work was supported by the Fraunhofer Internal Programs under Grant No. PREPARE 40-08394.
Code Availability
Appendix A Appendix / Supplemental Material
A.1 Influence Functions
In this part, we aim to formulate the key steps for influence functions, initially introduced in 1974 [11] to study robust estimators, and which were recently applied to neural networks [16; 17]. However, as already observed in [29], the functions typically encountered in neural networks require a more exhaustive analysis. Therefore, we will carefully redraw the results proposed in [16] and generalize them slightly for the application to PINNs. In particular, we will formulate influence functions in a manner that enables the study of the effect of various modifications of the problem (changes in the training set, in the loss weighting or in the PDE) on the prediction and any derived quantity. Furthermore, we will provide rigorous proofs for the statements made in Section 2.
We will start with a crucial lemma for influence functions, which was already noted in [6]. As remarked in [29], influence functions are usually stated for strictly convex functions [^1] Note, that for the application of influence functions with global minima (e.g. [16]), strict convexity is a necessary, but not a sufficient criterion. and relates to global minima. However, in practice, we will typically only reach saddle points [8], which is the reason why the following lemma is extended to stationary points.
Lemma 1.
For $\Theta\subseteq\mathbb{R}^{n}$ and $U\subseteq\mathbb{R}$ open, let $g:\Theta\times U\to\mathbb{R}$ be a twice continuously differentiable function for which there exists an $(\theta_{0},\epsilon_{0})\in\Theta\times U$ such that $\theta_{0}$ is a non-degenerate stationary point of the function $g(\cdot,\epsilon_{0})$, i.e. $\nabla_{\theta}g(\theta_{0},\epsilon_{0})=0$ and the Hessian $\mathcal{H}_{\theta}(\theta_{0},\epsilon_{0})\coloneq\nabla_{\theta}^{2}g(\theta_{0},\epsilon_{0})$ is invertible. Then there exists a non-empty, open set $U_{0}\subseteq U$ with $\epsilon_{0}\in U_{0}$ such that $g(\cdot,\epsilon)$ has a unique stationary point in the neighbourhood of $\theta_{0}$ for each $\epsilon\in U_{0}$ and we call the function describing those stationary points by $h:U_{0}\to\Theta$. This function $h$ is continuously differentiable with derivative
$\frac{\partial h(\epsilon)}{\partial\epsilon}=-\mathcal{H}_{\theta}(h(\epsilon),\epsilon)^{-1}\cdot\left.\nabla_{\theta}\frac{\partial}{\partial\epsilon}g(\theta,\epsilon)\right|_{\theta=h(\epsilon)}\quad\text{.}$ (7)
If $g$ is analytic in a neighbourhood of $(\theta_{0},\epsilon_{0})$, then we can choose $U_{0}$ such that $h$ is analytic as well. Furthermore, let $B\subseteq U$ be an open, connected set with $\epsilon_{0}\in B$, such that for every $\epsilon\in B$ there exists a non-degenerated stationary point of $g(\cdot,\epsilon)$. Then $h$ is extendable to $B$ and this function $h:B\to\Theta$ is continuous differentiable as well with Eq. 7.
Proof.
By $s:\Theta\times U\to\mathbb{R}^{n}$ we denote the gradient $s(\theta,\epsilon)\coloneq\nabla_{\theta}g(\theta,\epsilon)$, which is continuously differentiable by assumption. Since $(\theta_{0},\epsilon_{0})$ is a stationary point, this will be a root point of $s$. By the implicit function theorem (see e.g. [25, thm. 2]) we can conclude that there is an open set $U_{0}\subseteq U$ around $\epsilon_{0}$ such that there exists a unique function $h:U_{0}\to\Theta$ with $h(\epsilon_{0})=\theta_{0}$ and $s(h(\epsilon),\epsilon)=0$ for all $\epsilon\in U_{0}$. Therefore, for any $\epsilon\in U_{0}$, there is only one unique stationary point in the neighbourhood of $\theta_{0}$. Furthermore, $h$ is continuously differentiable, and its derivative is given by Eq. 7.
If $g$ is analytic, then the same holds for its gradient $s$. By the holomorphic implicit function theorem (see e.g. [9, sec. 7.6]) the function $h$ is also analytic.
If the assumptions of the lemma are satisfied for any $\epsilon\in B$, we can apply the lemma in all of these points. The function $h:B\to\Theta$ is continuous differentiable, since it is continuous differentiable in every point and Eq. 7 extends to all points in $B$. Moreover, $h(B)$ is connected, since $h$ is continuous and $B$ is connected by assumption. ∎
In a general sense, influence functions study the effect of a perturbation of a function, such as a loss function, through a functional, such as the argmin operator. Therefore, we will consider a functional $F:C(\Theta\times U)\times U\to\Theta$ and a continuous function $g:\Theta\times U\to\mathbb{R}^{m}$ with $\Theta\subseteq\mathbb{R}^{n}$ and a neighbourhood $U\subseteq\mathbb{R}$ containing the origin. We assume that $g_{0}(\theta)\coloneq g(\theta,0)$ describes the initial problem, whereas $g(\theta,\epsilon)$ for $\epsilon>0$ corresponds to a perturbed problem, e.g. changes in the training set. The influence function can then be formally expressed as
$\operatorname{Inf}F[g(\cdot,\cdot)]\coloneq\lim_{\epsilon\to 0}\frac{F[g(\cdot,0)]-F[g(\cdot,\epsilon)]}{\epsilon}$ (8)
provided that this limit exists.
In the case of stationary points (in particular for global minima) we obtain a much more convenient expression for the influence function by means of Lemma 1.
Theorem 1.
Let $F$ be the functional of stationary points [^2] To be a functional, $F$ have to map every function to one stationary point or one minimum, respectively. Hence, we might have to adjust e.g. the domain of the input functions to get unique stationary points. (alternatively the argmin operator) and let $g_{0}:\Theta\to\mathbb{R}$ and $\kappa:\Theta\to\mathbb{R}$ be twice continuously differentiable functions. Assume that $g_{0}$ has a non-degenerated, stationary point $\hat{\theta}=F[g_{0}]$ (a global minimum, respectively). Then the influence function is given by
$\operatorname{Inf}F[g_{0}+\epsilon\kappa]=\mathcal{H}_{\hat{\theta}}^{-1}\nabla_{\theta}\kappa(\hat{\theta})\quad\text{,}$ (9)
with the Hessian $\mathcal{H}_{\hat{\theta}}=\nabla_{\theta}^{2}g_{0}(\hat{\theta})$.
Usually we are interested in the effect of a change in the training process on the evaluation of a function $f$, such as the loss or the prediction. Therefore, before proving Theorem 1, we will add the following direct consequence of the definition (8).
Corollary 1.
Let $f:\Theta\to\mathbb{R}^{d}$ be a differentiable function. Then the influence function satisfies a chain rule
$\operatorname{Inf}f(F[g(\cdot,\cdot)])=\nabla_{\theta}f(\theta)\Big{|}_{\theta=F[g(\cdot,0)]}\cdot\operatorname{Inf}F[g(\cdot,\cdot)]\quad\text{.}$ (10)
Proof.
If the limit in (8) exists, it is nothing else than the derivative. Thus, by applying Lemma 1, the derivative exists and is given by
$\operatorname{Inf}F[g_{0}+\epsilon\kappa]=-\left.\frac{dF(g_{0}+\epsilon\kappa)}{d\epsilon}\right|_{\epsilon=0}=\mathcal{H}_{\hat{\theta}}^{-1}\nabla_{\theta}\kappa(\hat{\theta})\quad\text{.}$ (11)
The corollary directly follows from the chain rule for differentiation. ∎
Note that the left-hand side of Eq. 9 is the Gâteaux derivative of $F$ at $g_{0}$ in the direction of $\kappa$, which provides a convenient mathematical setting to study influence functions for linear perturbations. We refer to [12] for the study of those relations.
In practice, we often use infinitesimal changes expressed by $\operatorname{Inf}f(F[g(\cdot,\epsilon)])$ to approximate some finite change in $f(F[g(\cdot,\epsilon)])$ for an $\epsilon>0$. From this perspective, the influence function is the linear approximation term of a Taylor series
$f(F[g(\cdot,\epsilon)])=f(F[g(\cdot,0)])-\epsilon\cdot\operatorname{Inf}f(F[g(\cdot,\cdot)])+r(\epsilon)\quad\text{,}$ (12)
where the remainder term $r(\epsilon)$ vanishes with
$\lim_{\epsilon\to 0}\frac{r(\epsilon)}{\epsilon^{2}}=0$ (13)
If, moreover, the functions $f$ and $g$ are analytic, then by Lemma 1, also $f(F[g(\cdot,\epsilon)])$ is analytic and the Taylor series converges uniformly for every compact subset of its convergence region, i.e. for any compact subset of $\mathbb{R}$ we have a constant $C>0$ such that the remainder is bounded there by
$\left|r(\epsilon)\right|\leq C\epsilon^{2}\quad\text{.}$ (14)
In the remaining part of this section, we will study concrete settings for influence functions, which appears in particular for PINNs, and give convenient reformulations of the statement of Theorem 1.
Adding & removing training points:
Let $\mathcal{X}$ be a finite, non-empty set of training points and consider a finite set of points $\mathcal{X}^{+}$, which we would like to add to the training set, as well as a set of points $\mathcal{X}^{-}\subsetneq\mathcal{X}$, which we would like to remove from the training set. Moreover, let $L(\boldsymbol{x},\cdot):\Theta\to\mathbb{R}$ be a twice continuously differentiable loss function assigning a value to each (potential) training point of $\mathcal{X}\cup\mathcal{X}^{+}$. The empirical risk for our original dataset is given by $g_{0}(\theta)=\frac{1}{N}\sum_{\boldsymbol{x}\in\mathcal{X}}L(\boldsymbol{x};\theta)$ and let $\kappa(\theta)=\sum_{\boldsymbol{x}\in\mathcal{X}^{+}}L(\boldsymbol{x};\theta)-\sum_{\boldsymbol{x}\in\mathcal{X}^{-}}L(\boldsymbol{x};\theta)$ be the perturbation in risk corresponding to the change in the training set. $N$ denotes the size of training points in $\mathcal{X}$. Assume that there exists a stationary point [^3] Thus, in practice, we assume that our training has progressed to the stage where it has reached a stationary point in the loss landscape. (a global minimum, respectively) $\hat{\theta}=F[g_{0}]$ inside $\Theta$ and let $F$ be the functional which maps to those stationary points. In addition, let $f(\boldsymbol{z},\cdot):\Theta\to\mathbb{R}^{m}$ be an arbitrary continuously differentiable function (e.g. the loss or the prediction), possibly with a further dependency expressed by a variable $\boldsymbol{z}$ (e.g. a test point).
According to Theorem 1, we can then approximate counterfactual effect of training on a perturbed training set $\mathcal{X}^{\prime}=(\mathcal{X}\cup\mathcal{X}^{+})\setminus\mathcal{X}^{-}$ by
$f(\boldsymbol{z};\hat{\theta})-f(\boldsymbol{z};\hat{\theta}^{\prime})=\frac{1}{N}\operatorname{Inf}_{f(\boldsymbol{z};\hat{\theta})}(\mathcal{X}^{+},\mathcal{X}^{-})+r\!\left(N^{-2}\right)\quad\text{,}$ (15)
where we introduce the shorthand notation
$\operatorname{Inf}_{f(\boldsymbol{z};\hat{\theta})}(\mathcal{X}^{+},\mathcal{X}^{-})$ :-$\coloneq\operatorname{Inf}f(\boldsymbol{z};F[g_{0}+\epsilon\kappa]):$
$=\nabla_{\theta}f(\boldsymbol{z};\hat{\theta})\cdot\mathcal{H}_{\hat{\theta}}^{-1}\cdot\nabla_{\theta}\left(\sum_{\boldsymbol{x}\in\mathcal{X}^{+}}L(\boldsymbol{x};\hat{\theta})-\sum_{\boldsymbol{x}\in\mathcal{X}^{-}}L(\boldsymbol{x};\hat{\theta})\right)\quad\text{,}$ (16)
which is simply the signed sum of the influences of the individual training points due to the linearity of Influence Functions.
Removing a complete loss term:
Moreover, we can also remove a complete loss term and study the influence of this operation to the training. For this, consider a total loss as a sum of $k$ loss terms
$g_{0}(\theta)=\mathcal{L}(\mathcal{X};\theta)=\sum_{j=1}^{k}\mathcal{L}_{j}(\mathcal{X};\theta)=\frac{1}{N}\sum_{\boldsymbol{x}\in\mathcal{X}}\sum_{j=1}^{k}L_{j}(\boldsymbol{x};\theta)$ (17)
[…]
with
$\operatorname{Inf}f(\boldsymbol{z};F[g_{0}+N^{-1}\kappa])=-\nabla_{\theta}f(\boldsymbol{z};\hat{\theta})\cdot\mathcal{H}_{\hat{\theta}}^{-1}\cdot\nabla_{\theta}\sum_{\boldsymbol{x}\in\mathcal{X}}L_{i}(\boldsymbol{x};\hat{\theta})\quad\text{.}$ (19)
Influence of loss weights:
In analogous fashion, we can easily extend this framework to study the effect of upweighting or downweighting individual loss terms. This is in particular interesting for PINNs because a wrong weighting of the loss terms is known to reduce the accuracy of PINNs drastically [31].
Thus, assume that for some weights $w_{j}>0$ we consider a total loss in the form of
$g_{0}(\theta)=\mathcal{L}(\mathcal{X};\theta)=\sum_{j=1}^{k}w_{j}\mathcal{L}_{j}(\mathcal{X};\theta)$ (20)
where we now perturb the weight $w_{i}$ with $\kappa(\theta)=w_{i}\mathcal{L}_{i}(\mathcal{X};\theta)$. The influence of changing the weight $w_{i}$ is then given by
$\operatorname{Inf}f(\boldsymbol{z};F[g_{0}+\epsilon\kappa])=w_{i}\nabla_{\theta}f(\boldsymbol{z};\hat{\theta})\cdot\mathcal{H}_{\hat{\theta}}^{-1}\cdot\nabla_{\theta}\mathcal{L}_{i}(\mathcal{X},\hat{\theta})\quad\text{.}$ (21)
In comparison to the two previous discussed cases, we can also consider infinitesimal changes of the weight so that we can omit a discussion of the error bounds.
Influence of parameters in the PDE:
As a last option to use influence functions for PINNs, we want to discuss the impact of parameters in the PDE. Therefore, we will consider a basic objective of the form
$g(\theta,\rho)=\frac{1}{N}\sum_{\boldsymbol{x}\in\mathcal{X}}L(\boldsymbol{x};\rho;\theta)$ (22)
with some additional parameter $\rho\in\mathbb{R}$, and we will assume that $\mathcal{L}$ is also continuously differentiable w.r.t. $\rho$. To apply Lemma 1, we will assume, that there is an optimal or stationary parameter $\hat{\theta}=F[g(\cdot,\rho_{0})]$ at a certain value of the parameter $\rho_{0}$ we are interested in. Then, the influence of the parameter $\rho$ to the optimal parameters $\hat{\theta}$ at a certain point $\rho_{0}$ is given by
$\operatorname{Inf}F[g]$ $=\lim_{\epsilon\to 0}\frac{F[g(\theta,\rho_{0})]-F[g(\theta,\rho_{0}+\epsilon)]}{\epsilon}=-\left.\frac{dF[g(\theta,\rho)]}{d\rho}\right|_{\rho=\rho_{0}}$
$=\mathcal{H}_{\theta}^{-1}(\hat{\theta},\rho_{0})\cdot\nabla_{\theta}\frac{\partial}{\partial\rho}g(\hat{\theta},\rho_{0})\quad\text{.}$ (23)
A.2 Navier-Stokes Parametrization
The rectangular cavity characterizing the 2D Navier-Stokes problem formulation is given in Table 2.
Table 2: Geometric Parameters of the Cavity
Parameter Value [$\mathrm{m}$]
$x_{\min}$ 0
$x_{\max}$ 2.2
$y_{\min}$ 0
$y_{\max}$ 0.41
Cylinder radius 0.05
Cylinder position $(x_{c},y_{c})$ (0.2,0.2)
The parabolic inflow is defined as
$\mathbf{u}(x_{\min},y)$ $=\begin{pmatrix}U_{\text{max}}y\frac{y_{\max}-y}{y_{\max}^{2}}\\ 0\end{pmatrix},$ (24)
[…]
A.3 Supplemental Figures
Figure 4 is supplementary to Figure 1 and shows the model predictions additionally for $|\!|\boldsymbol{u}|\!|$ and for $\phi_{\text{broken}}$, $\phi_{\text{bad}}$ as well as target values precomputed with the finite element method (FEM) program FEniCS sourced from [33].
Figure 5 shows the absolute errors between model predictions and the target values depicted in Figure 4(d). Note that $\phi_{\text{broken}}$ is evaluated against correct target values.
Figure 6 shows the influence heatmap produced by $\phi_{\text{broken}}$ and $\phi_{\text{bad}}$ for the same training point as used in Figure 2.
(a) $\phi_{\text{good}}$
(b) $\phi_{\text{broken}}$
(c) $\phi_{\text{bad}}$
(d) Target values computed with FEniCS.
Figure 4: Trained models and target values with respective predictions for $u_{1}$, $u_{2}$, $p$, $|\!|\boldsymbol{u}|\!|$ (top to bottom).
(a) $\phi_{\text{good}}$
(b) $\phi_{\text{broken}}$
(c) $\phi_{\text{bad}}$
Figure 5: Absolute errors between respective model predictions for each output dimension w.r.t. precomputed target values.
(a) $\phi_{\text{broken}}$
(b) $\phi_{\text{bad}}$
Figure 6: Influence heatmaps on the prediction of $u_{1}$ of a single training point close to the cylinder (bottom).